# EntroLLM: Entropy Encoded Weight Compression for Efficient Large Language Model Inference on Edge Devices

### v3

@@ IV Experiments
We evaluated our proposed compression scheme on three edge-based LLMs: smolLM-1.7B-Instruct (1.7 billion parameters) [9], phi3-mini-4k-Instruct (3.8 billion parameters) [10] and mistral-7B-Instruct (7 billion parameters) [4]. Table I shows their baseline fp16 sizes and subsequent sizes after quantization and Huffman encoding. Figure 4 visually confirms the Gaussian distribution of the quantized weights.
IV-A Storage Reduction
@@
TABLE II: Latency breakdown for the phi3-mini-4k model on an NVIDIA Jetson P3450 across different quantization formats (uint8 and uint4) with and without Huffman compression.
Storage reduction is achieved through a two-step process. First, the models undergo quantization. For example, the phi3-mini-4k-Instruct model’s size is reduced from $7.2$ GB to $3.6$ GB and $1.8$ GB when quantized from fp16 to uint8 and uint4, respectively. This initial step significantly reduces the memory footprint of the models.
Following quantization, Huffman compression is applied. This further reduces the effective bit lengths from $8$ and $4$ bits (to $5.58$ and $1.39$ bits, respectively, in the case of phi3-mini-4k-Instruct). Table I provides additional details for the bit widths for the other two models, smolLM-1.7B-Instruct and mistral-7B-Instruct.
The transition from $8$-bit to $4$-bit quantization increases the skewness of the distribution. This shift happens because reducing bit-width compresses the number of discrete values available to represent parameters. In the $8$ bit quantization, the $256$ symbols allow for a more granular weight representation, while the $4$-bit quantization reduces this to only $16$ symbols. This reduction causes a “bucketing” effect, where multiple nearby values are mapped to the same symbol, reducing the randomness or entropy of the distribution. The central buckets, which contain the most frequent values, dominate the quantized distribution, especially in the 4-bit quantized model.
IV-B Performance
@@
We tested our compressed models on a range of tasks to evaluate their performance under low-precision quantization. Specifically, we considered text generation on the Wikitext2 benchmark, common sense reasoning on the 5-shot HellaSwag benchmark, and arithmetic reasoning on the 8-shot GSM8k CoT benchmark. Table I summarizes the results for three model families—smolLM-1.7B-Instruct, phi3-mini-4k-Instruct mistral-7B-Instruct – evaluated across fp16, uint8, and uint4 precision levels.
Across all models and benchmarks, the uint8 and uint4 variants exhibit only a modest performance degradation compared to the fp16 baseline. For example, on Wikitext2, perplexity remains largely stable – phi3-mini-4k-Instruct shifts from $9.03$ to $10.10$, and mistral-7B-Instruct degrades slightly from $8.17$ to $8.29$, despite operating at just $1.39$ and $1.62$ effective bits, respectively, in the uint4 format. These results highlight the resilience of the language modeling performance of our LLMs to aggressive quantization.
For HellaSwag, accuracy remains largely consistent across bit widths, especially for larger models. The model phi3-mini-4k-Instruct, for example, drops only $1.2$ percentage points between fp16 and uint4, maintaining over $81\%$ accuracy. The smaller smolLM-1.7B-Instruct shows a slightly larger relative drop, suggesting that model size plays a role in quantization robustness.
For GSM8k CoT, which is more sensitive to precision due to its dependence on intermediate reasoning steps, we observe a larger performance gap. The phi3-mini-4k-Instruct model drops from $77.37\%$ to $70.58\%$ in accuracy when moving from fp16 to uint4, while mistral-Instruct decreases from 52.2% to 45.36%. However, given the compression achieved, this drop may still be acceptable for many deployment scenarios.
IV-C Hardware Efficiency & Parallel Decoding Performance
Table II presents latency measurements for various stages of inference using the phi3-mini-4k-Instruct model, which contains approximately $3.8$ billion parameters. The evaluation was performed on an NVIDIA Jetson P3450, where we parallelized the decoding operation across four CPU threads.
NVIDIA Jetson P3450 features a quad-core ARM Cortex-A57 CPU running at $1.43$ GHz, $4$ GB LPDDR4 memory with $25.6$ GB/s bandwidth, $32$ KB L1 data cache per core, $48$ KB L1 instruction cache per core, $2$ MB shared L2 cache, and a $128$-core Maxwell GPU. The Cortex-A57 cores support NEON SIMD instructions [28], enabling efficient bit-level parallelism via packing in C++ and significantly accelerating low-precision Huffman decoding operations like those used in our uint4 and uint8 models.
Our results show that parallel decoding is highly efficient, especially under aggressive quantization. For the uint4 model, decoding takes only $1.66$ seconds, a small number compared to the $9.69$ seconds required for pre-fill and $9.75$ seconds required to generate the first output token. Even with uint8, parallel Huffman decoding is completed in $6.66$ seconds, which remains a small fraction of the overall inference time. Crucially, Huffman decoding is performed only once per sequence. As a result, its cost is amortized over the full output length, making the overhead negligible in the overall inference pipeline. This is especially significant in benchmarks like Wikitext2, HellaSwag, and GSM8k, where each input may result in dozens or hundreds of tokens. The ability to decode all quantized weights efficiently upfront enables fast, low-power inference at the edge, showing that our compressed models are not only memory and bandwidth-efficient but also latency-friendly for real-world deployments.
IV-D Latency Reduction from Huffman Coding
Table II demonstrates the performance benefits of Huffman coding on the phi3-mini-4k model running on an NVIDIA Jetson P3450. The data reveal significant throughput and latency improvements when applying Huffman coding on quantized weights of $8$-bit and $4$-bit. For the $8$-bit (uint8) format, Huffman coding reduces the effective bit width to $5.58$ bits, yielding a $14.5\%$ reduction in pre-fill time (from $27.10$ to $23.17$ seconds) and a $31.9\%$ decrease in token generation latency (from $0.083$ sec to $0.063$ sec). Even more impressive, for $4$-bit (uint4) weights, Huffman compression achieves an effective bit-width of just $1.39$ bits, resulting in a $13.3\%$ pre-fill time reduction and a remarkable $146.6\%$ improvement in token generation latency (from $0.062$ sec to $0.025$ sec). These results demonstrate that reducing the memory bandwidth requirements through Huffman coding has a dramatic impact on inference performance, particularly for token generation. Performance gains from Huffman coding are particularly pronounced in the LLM decoding phase, which is primarily memory-bandwidth limited [29, 30]. With reduced bit-width, less data needs to be transferred from memory, directly improving the latency to generate an output token. For example, when $8$-bit weights are compressed to $5.58$ bits, theoretically $30\%$ fewer data needs to be moved, approaching a 1.43×1.43$\times$1.43 × potential speedup. Our measured 1.32×1.32$\times$1.32 × improvement $\left(\frac{15.93}{12.08}=1.32\times\right)$ closely approaches this theoretical limit, with the difference attributable to bit-packing overheads. The improvements in pre-fill latency are more modest because this phase is compute-dominated rather than memory-bandwidth limited [31]. Although our approach reduces memory access with lower bit-width, it does not reduce the computational complexity since Huffman encoded weights are decoded back to their original precision ($4$ or $8$ bits) before computation. Therefore, the primary benefit during pre-fill comes from reduced data transfer from on-chip memory and IO, not from the compute operations themselves. As discussed in Section IV-C, Huffman decoding is performed once per sequence and introduces negligible overhead, especially when amortized across long outputs common in LLM workloads.
The implementation challenges of non-standard bit formats are effectively managed through optimized CUDA kernels that efficiently pack and unpack these fractional bit-width values [32]. While current Ampere GPU Tensor Cores natively support int8 and int4 operations but not arbitrary bit-widths as obtained during Huffman coding, our approach minimizes overhead by keeping the computational precision unchanged and focusing on memory transfer optimization.
V Conclusions & Discussions
In this work, we have developed a highly efficient method for compressing and deploying large language models (LLMs) on resource-constrained edge devices. Our approach combines low-bit quantization, entropy coding via Huffman compression, and a parallelized decoding strategy to achieve significant storage and bandwidth reductions – without compromising model accuracy or inference throughput. Specifically, our mixed symmetric and asymmetric quantization strategy effectively preserves the natural weight distribution across layers, optimizing the compressibility of the quantized tensors. This enables Huffman coding to reduce the average bit-width to as low as $1.39$ bits for $4$-bit weights, yielding dramatic improvements in memory efficiency. To address the inherent latency challenges of entropy decoding, we introduce a lightweight address space segmentation scheme that enables parallel Huffman decoding, ensuring that the decompression overhead remains negligible and scalable even on devices with limited compute, such as the NVIDIA Jetson P3450. Experimental results validate the efficiency of our method in the real world: we observe substantial improvements in token generation latency, especially in memory-bound inference phases such as LLM decoding. In general, our pipeline facilitates real-time, low-power inference for quantized LLMs on embedded platforms, making it suitable for on-device natural language processing in applications such as robotics, Internet-of-things, autonomous vehicles, and edge AI assistants. Our future work includes exploring adaptive entropy coding and hardware-aware optimizations to further reduce overhead and improve deployment efficiency.

### v4

@@ 4 Experiments
We evaluated our proposed compression scheme on three edge-based LLMs: smolLM-1.7B-Instruct (1.7 billion parameters) [2], phi3-mini-4k-Instruct (3.8 billion parameters) [1] and mistral-7B-Instruct (7 billion parameters) [18]. Our code and models are available online [26]. Table 1 shows their baseline fp16 sizes and subsequent sizes after quantization and Huffman encoding. Figure 4 visually confirms the Gaussian distribution of the quantized weights.
[…]
@@
Table 3: Latency breakdown for the phi3-mini-4k model on an NVIDIA Jetson P3450 across different quantization formats (uint8 and uint4) with and without Huffman compression.
@@
Performance: We tested our compressed models on a range of tasks to evaluate their performance under low-precision quantization. Specifically, we considered text generation on the Wikitext2 benchmark, common sense reasoning on the 5-shot HellaSwag benchmark, and arithmetic reasoning on the 8-shot GSM8k CoT benchmark. Table 1 summarizes the results for three model families—smolLM-1.7B-Instruct, phi3-mini-4k-Instruct mistral-7B-Instruct – evaluated across fp16, uint8, and uint4 precision levels.
Hardware Efficiency & Parallel Decoding: Table 3 shows latency measurements for the phi3-mini-4k-Instruct model (3.83.8B parameters) evaluated on an NVIDIA Jetson P3450 with four CPU threads for parallel decoding. The device features a quad-core ARM Cortex-A57 CPU ($1.43$ GHz), $4$ GB LPDDR4 memory (25.6 GB/s), and a 128-core Maxwell GPU. NEON SIMD support [15] enables efficient bit-level parallelism, accelerating low-precision Huffman decoding for uint4 and uint8 models. Parallel decoding is highly efficient, especially for uint4, where decoding takes only $1.66$ s versus $9.69$ s for pre-fill and $9.75$ s for the first output token. Even for uint8, decoding completes in $6.66$ s, a small fraction of total inference time. Since decoding occurs only once per sequence, its cost is amortized over many tokens, making the overhead negligible. This enables fast, low-power, and latency-friendly inference on edge benchmarks like Wikitext2, HellaSwag, and GSM8k, demonstrating the practicality of our compressed models for edge deployments.
Latency Reduction from Huffman Coding: Table 3 highlights the performance gains from applying Huffman coding to quantized phi3-mini-4k weights on the NVIDIA Jetson P3450. For uint8, Huffman coding reduces the effective bit-width to $5.58$ bits, cutting pre-fill time by $14.5\%$ and token generation latency by $31.9\%$. For uint4, it compresses to $1.39$ bits, yielding $13.3\%$ faster pre-fill and a $146.6\%$ improvement in token generation. These gains stem from reduced memory transfers during the memory-bandwidth-limited decoding phase [24, 27]. When 8-bit weights compress to $5.58$ bits, $30\%$ less data is moved, yielding a theoretical $1.43\times$ speedup, closely matched by our measured $1.32\times$. Pre-fill improvements are smaller since this phase is compute-bound [20]. Huffman decoding is done once per sequence, adding negligible overhead when amortized over many tokens. Although GPUs lack native support for fractional bit-widths, optimized CUDA kernels efficiently pack/unpack values [10], keeping compute precision fixed and focusing on memory transfer optimization.
5 Conclusions
We present EntroLLM, a framework combining mixed quantization, Huffman compression, and parallel decoding to enable efficient LLM deployment on edge devices. Our method reduces average bit-width to 1.39 for 4-bit weights, improving downstream entropy coding by $7\times$$11.3\times$ over state-of-the-art techniques. Parallel decoding keeps decompression overhead negligible, even on low-power hardware like the NVIDIA Jetson P3450. Experiments show significant latency and storage gains, enabling real-time, low-power inference for applications such as robotics and IoT. Future work will explore adaptive entropy coding and hardware-aware optimizations.